The vertices of a triangle are A(-1,-7) B(5,1) and c(1,4). The equation of the bisector of angle Abc is ??

Grade:11

1 Answers

Dheeraj
5 years ago
First find the equation of  intersecting sides
Let it be a1x+b1y=c1
  a2x+b2y=c2
 
 
 
Now use formula
(a1x+b1y+c)/√a1^2+b1^2=+-(a2x +b2y+c)/√a2^2+b2^2
 You will find equation of bisector
Thankyou